OpenStudy (anonymous):

Simplify the expression. (k^2)^4

OpenStudy (campbell_st):

this uses the power of a power law \[(x^a)^b+ x^{a \times b}\] applying this to your problem \[(k^2)^4 = k^{2 \times4} = k^8\]